# Token-based Audio Inpainting via Discrete Diffusion

### v3

@@ Abstract
Audio inpainting seeks to restore missing segments in degraded recordings. Previous diffusion-based methods exhibit impaired performance when the missing region is large. We introduce the first approach that applies discrete diffusion over tokenized music representations from a pre-trained audio tokenizer, enabling stable and semantically coherent restoration of long gaps. Our method further incorporates two training approaches: a derivative-based regularization loss that enforces smooth temporal dynamics, and a span-based absorbing transition that provides structured corruption during diffusion. Experiments on the MusicNet and MAESTRO datasets with gaps up to 750 ms show that our approach consistently outperforms strong baselines across range of gap lengths, for gaps of 150 ms and above. This work advances musical audio restoration and introduces new directions for discrete diffusion model training. Audio examples of our proposed method can be found at https://iftach21.github.io/.
@@ 1 Introduction
Audio inpainting refers to the task of reconstructing missing or corrupted segments of an audio signal (adler2011audio, ). It is a fundamental inverse problem in audio processing, with applications ranging from restoring damaged recordings and removing artifacts, to filling in gaps caused by data loss in transmission or editing operations (moliner2023diffusion, ). Traditional approaches to this problem often relied on signal modeling techniques such as autoregressive models (adler2011audio, ), sparse representations (marafioti2020gacela, ), or linear predictive coding (adler2011audio, ). While effective under certain assumptions (e.g., local stationarity), these methods typically perform well only on short gaps and may struggle with long-range dependencies or semantic coherence (moliner2023diffusion, ; zhou2019vision, ).
In recent years, deep generative models have significantly advanced the state of audio inpainting. Models such as Variational Autoencoders (VAEs) (marafioti2019context, ) and diffusion probabilistic models (kong2020diffwave, ; liu2023maid, ; ho2020denoising, ) have demonstrated the ability to learn expressive priors directly from large-scale audio datasets. Notably, diffusion models have emerged as particularly powerful tools for solving ill-posed inverse problems due to their iterative denoising process and strong generative capacity. Approaches such as DiffWave (kong2020diffwave, ) apply diffusion directly on waveform samples, while others such as MAID (liu2023maid, ) and CQT-Diff+ (moliner2023diffusion, ) operate in the continuous time-frequency domain using spectrograms or Constant-Q Transform (CQT) representations (moliner2023diffusion, ; zhou2019vision, ).
@@
In this work, we propose a novel approach: Audio Inpainting via Discrete Diffusion (AIDD). Unlike prior methods that operate in the continuous domain (kong2020diffwave, ; liu2023maid, ; moliner2023diffusion, ; zhou2019vision, ), our method applies the diffusion process to a discrete latent space. Specifically, we employ the WavTokenizer (ji2024wavtokenizer, ) to quantize audio signals into compact sequences of discrete tokens, and perform the diffusion process entirely in this categorical space. This formulation allows the model to capture high-level semantic structures in audio, while avoiding the challenges of modeling raw waveforms or spectrograms directly. To address these challenges, we introduce a discrete diffusion framework for audio inpainting. Our contributions are threefold, advancing audio inpainting with discrete diffusion model:
@@
  • Performance. We present comparable or state-of-the-art results for audio inpainting on MusicNet (dataset, ) and MAESTRO (MAESTRO, ) dataset, for gaps ranging from 150 - 750 ms. Showing consistent results for both large and small datasets.
@@ 2.1 Methods for Audio Inpainting
The term “audio inpainting” was introduced in (adler2011audio, ) to describe the restoration of missing or corrupted audio segments, although related notions had appeared under names such as audio interpolation (marafioti2020gacela, ), audio extrapolation (lieb2018audio, ), missing sample reconstruction, waveform substitution (adler2011audio, ), and imputation (lieb2018audio, ).
Classical methods mainly targeted short gaps (<100<100 ms), often assuming local stationarity. Auto-regressive models (adler2011audio, ) predicted missing samples from past ones but degrade on longer gaps. Sparse reconstruction in time–frequency domains like STFT or Gabor (lieb2018audio, ; adler2011audio, ) exploited algorithms such as OMP (tropp2007signal, ), later improved by adaptive dictionaries (e.g., learned Gabor atoms (taubock2020dictionary, )). NMF-based methods (mokry2023algorithms, ) filled spectrogram gaps via low-rank factorization, with probabilistic variants improving robustness. Other structured priors include sinusoidal modeling for harmonic signals (adler2011audio, ) and graph-based regularization: (perraudin2018inpainting, ) used a graph Laplacian on self-similarity matrices to borrow content from uncorrupted regions.
While effective for very short gaps, these models fail for longer ones due to violated assumptions of stationarity or sparsity. Deep learning approaches were introduced to overcome this challenge. CNNs first inpainted spectrograms, with (marafioti2019context, ) using a U-Net style context encoder for tens-of-milliseconds gaps. GANs further improved realism: (ebner2020audio, ) proposed a Wasserstein GAN with multi-context conditioning for gaps up to 500 ms, while (marafioti2020gacela, ) extended this with GACELA, a multi-scale GAN handling 1–1.5 s gaps. These data-driven models substantially outperform classical approaches, especially on long gaps with complex temporal or musical structure. More recently, A score-aware GAN framework (aironi2023score, ) further improved music inpainting by incorporating symbolic losses, outperforming GACELA on medium gaps.
@@ 2.2 Audio Inpainting using Diffusion Models
Diffusion models, or score-based generative models, have recently transformed vision and audio generation. They define a forward noising process and learn its reversal (e.g., DDPMs (ho2020denoising, ), score-SDEs (song2020score, )). In audio, they support waveform and spectrogram generation conditioned on context. For instance, DiffWave (kong2020diffwave, ) synthesizes raw waveforms for vocoding, while other works apply diffusion in latent or time–frequency domains (STFT, CQT) to exploit structure (moliner2023solving, ; moliner2023diffusion, ). Most use continuous diffusion, though discrete approaches exist: DiffSound (yang2023diffsound, ) applies discrete diffusion to quantized spectrogram tokens. Conceptually, continuous diffusion perturbs real values, yielding smooth interpolations, whereas discrete diffusion masks/replaces categorical tokens, enabling multimodal reconstructions but with possible quantization artifacts (ho2020denoising, ; song2020score, ; yang2023diffsound, ).
Diffusion has recently been adapted to inpainting. (moliner2023diffusion, ) propose a zero-shot unconditional model on CQT spectrograms, with posterior sampling to enforce data consistency (kawar2022denoising, ; chung2022diffusion, ). This handles arbitrary gaps, matching baselines on short ones and surpassing them up to 300 ms. Similarly, CQT-Diff (moliner2023solving, ) uses an invertible CQT front-end for inpainting, declipping, and bandwidth extension, achieving state-of-the-art restoration. MAID (liu2023maid, ), instead, trains a conditional diffusion model for music inpainting guided by context and auxiliary cues (e.g., pitch), enabling longer gaps but requiring task-specific supervision. (aironi2023score, ) further propose score-aware conditioning, aligning the score function with masked regions for musically coherent results.
Architectures differ in operating domain (raw waveform vs. spectrogram/latent space) and conditioning. Zero-shot approaches (moliner2023diffusion, ) enforce context consistency during inference, while conditional models integrate masked audio or side information during training. Masking usually involves contiguous spans: Moliner et al. test gaps from 25–300 ms, while GAN baselines address seconds-long gaps (marafioti2020gacela, ; ebner2020audio, ). Diffusion models degrade gracefully with gap length: all interpolate short gaps well, but diffusion’s generative prior yields more plausible completions beyond 100 ms compared to sparsity- or AR-based methods.
In summary, diffusion-based inpainting combines strong generative priors with flexible conditioning. Open questions remain on continuous vs. discrete diffusion (yang2023diffsound, ) and conditioning strategies (posterior sampling vs. trained conditional networks).
@@ 3.1 Discrete Diffusion Models
Continuous Diffusion Models (CDMs) (ho2020denoising, ; song2020score, ) have achieved remarkable success across various generative modeling tasks, particularly in the image domain (dhariwal2021diffusion, ). The high-dimensional nature of raw audio data adds to the difficulty of directly modeling it within the diffusion framework. An effective strategy to address this issue is to first compress continuous audio signals into compact, discrete representations, such as quantized tokens derived from a learned codebook (defossez2022high, ). DDMs, which operate in a token space, have recently demonstrated strong performance in the domain of natural language generation (lou2310discrete, ; nie2025large, ; sahoo2024simple, ). Similarly to CDM, DDMs comprise a forward noising process that progressively corrupts an initial sample $x_{0}$ into a maximally noisy (fully masked) version $x_{T}$, and a reverse denoising process trained to reconstruct $x_{0}$ from $x_{T}$.
@@
Existing work (lou2310discrete, ) shows that the forward diffusion in discrete space can be formulated using a token-level transition rate matrix $Q_{\text{tok}}\in\mathbb{R}^{n\times n}$, where $n$ is the vocabulary size. The forward transition probability of a single token $x_{0}$ at time $t$ is then given by
@@
The reverse process is described by a new reverse transition matrix (sun2022score, ; kelly1981reversibility, ) which is defined as $\bar{Q}_{t}(y,x)=\tfrac{p_{t}(y)}{p_{t}(x)}Q_{t}(x,y)$ and $\bar{Q}_{t}(x,x)=-\sum_{y\neq x}\bar{Q}_{t}(y,x)$.
To simulate the reverse process, it is sufficient to estimate the ratio $\frac{p_{t}(y)}{p_{t}(x)}$, referred to as the concrete-score (lou2310discrete, ; meng2022concrete, ), which can be approximated by a neural score function: $s_{\theta}(x,t)\approx\left[\tfrac{p_{t}(y)}{p_{t}(x)}\right]_{y\in\mathcal{X},\,y\neq x}$.
@@
To train this score function, recent work has proposed the Diffusion Weighted Denoising Score Entropy (DWDSE) objective (lou2310discrete, ), which is defined as:
@@ 4 Audio Inpainting using Discrete Diffusion Model
To the best of our knowledge, our method Audio Inpainting using Discrete Diffusion (AIDD) is the first to leverage a discrete diffusion model for audio inpainting. AIDD comprises three main components: (1) audio tokenization using WavTokenizer (ji2024wavtokenizer, ), (2) generative modeling using a DDM (lou2310discrete, ), and (3) waveform reconstruction via token decoding (ji2024wavtokenizer, ). Furthermore, we propose a new training derivative-based loss, and a span-based masking method, extending the approach of (lou2310discrete, ) to more effectively address the audio inpainting task.
@@ 4.1 The Proposed Architecture
AIDD consists of WavTokenizer (ji2024wavtokenizer, ) that transforms high-dimensional raw audio signals into compact sequences of discrete tokens and reconstructs it back to raw audio, and a Diffusion Transformer (DiT) architecture (peebles2023scalable, ) which learns to predict masked tokens through the reverse diffusion process. An overview of the proposed framework is illustrated in Figure 1.
Figure 1: Our method operates on audio signals with missing (silent) segments. During inference, the input waveform, containing a silence gap is processed by the WavTokenizer encoder, which converts the audio into a discrete sequence of tokens. Next, a DiT performs inpainting by iteratively predicting the masked tokens, resulting in reconstructed token sequence. Finally, the reconstructed tokens are passed through the WavTokenizer’s decoder to synthesize the output audio waveform in the masked part. During training, token sequences are corrupted with span-based masking at randomly sampled timesteps, and the DiT is optimized to predict the concrete score using the DWDSE objective, complemented by the derivative-based loss.
@@
The first step in our method is to convert the raw audio waveform into discrete audio tokens that can be efficiently processed and trained by the DDM. We utilize a pre-trained audio tokenizer WavTokenizer (ji2024wavtokenizer, ), which compresses high-resolution audio into a compact sequence of discrete tokens using a single quantizer. Despite its extreme compression, WavTokenizer preserves high reconstruction fidelity and rich semantic content. Unlike conventional audio inpainting frameworks that operate directly on the audio waveform or the spectrogram (moliner2023diffusion, ), audio tokenization allows us to formulate audio inpainting as a discrete sequence completion task.
@@
At the core of our method is a DiT architecture (peebles2023scalable, ), which integrates time conditioning into a standard encoder-only transformer (vaswani2017attention, ; devlin2019bert, ) and, following the work of (lou2310discrete, ), we incorporate rotary positional encoding (su2024roformer, ). To operate effectively in the discrete token space, we use the denoising score entropy formulation, guiding the reverse diffusion dynamics as seen in Eq. 3.

### v4

@@ Abstract
Audio inpainting seeks to restore missing segments in degraded recordings. Previous diffusion-based methods exhibit impaired performance when the missing region is large. We introduce the first approach that applies discrete diffusion over tokenized music representations from a pre-trained audio tokenizer, enabling stable and semantically coherent restoration of long gaps. Our method further incorporates two training approaches: a derivative-based regularization loss that enforces smooth temporal dynamics, and a span-based absorbing transition that provides structured corruption during diffusion. Experiments on the MusicNet and MAESTRO datasets with gaps up to 750 ms show that our approach consistently outperforms strong baselines across range of gap lengths, for gaps of 150 ms and above. This work advances musical audio restoration and introduces new directions for discrete diffusion model training. Visit our project page for examples and code.
@@ 1 Introduction
Audio inpainting refers to the task of reconstructing missing or corrupted segments of an audio signal (Adler et al., 2011). It is a fundamental inverse problem in audio processing, with applications ranging from restoring damaged recordings and removing artifacts, to filling in gaps caused by data loss in transmission or editing operations (Moliner and Välimäki, 2023). Traditional approaches to this problem often relied on signal modeling techniques such as autoregressive models (Adler et al., 2011), sparse representations (Marafioti et al., 2020), or linear predictive coding (Adler et al., 2011). While effective under certain assumptions (e.g., local stationarity), these methods typically perform well only on short gaps and may struggle with long-range dependencies or semantic coherence (Moliner and Välimäki, 2023; Zhou et al., 2019).
In recent years, deep generative models have significantly advanced the state of audio inpainting. Models such as Variational Autoencoders (VAEs) (Marafioti et al., 2019) and diffusion probabilistic models (Kong et al., 2020; Liu et al., 2023; Ho et al., 2020) have demonstrated the ability to learn expressive priors directly from large-scale audio datasets. Notably, diffusion models have emerged as particularly powerful tools for solving ill-posed inverse problems due to their iterative denoising process and strong generative capacity. Approaches such as DiffWave (Kong et al., 2020) apply diffusion directly on waveform samples, while others such as MAID (Liu et al., 2023) and CQT-Diff+ (Moliner and Välimäki, 2023) operate in the continuous time-frequency domain using spectrograms or Constant-Q Transform (CQT) representations (Moliner and Välimäki, 2023; Zhou et al., 2019).
@@
In this work, we propose a novel approach: Audio Inpainting via Discrete Diffusion (AIDD  [^1] https://github.com/iftachShoham/AIDD). Unlike prior methods that operate in the continuous domain (Kong et al., 2020; Liu et al., 2023; Moliner and Välimäki, 2023; Zhou et al., 2019), our method applies the diffusion process to a discrete latent space. Specifically, we employ the WavTokenizer (Ji et al., 2024) to quantize audio signals into compact sequences of discrete tokens, and perform the diffusion process entirely in this categorical space. This formulation allows the model to capture high-level semantic structures in audio, while avoiding the challenges of modeling raw waveforms or spectrograms directly. To address these challenges, we introduce a discrete diffusion framework for audio inpainting. Our contributions are threefold, advancing audio inpainting with discrete diffusion model:
@@
  • Performance. We present comparable or state-of-the-art results for audio inpainting on MusicNet (Thickstun et al., 2016) and MAESTRO (Hawthorne et al., 2019) datasets, for gaps ranging from 150 - 750 ms, for audio segments containing either a single gap or multiple gaps. We show consistent results for both large and small datasets in both objective and subjective evaluations.
@@ 2.1 Methods for Audio Inpainting
The term “audio inpainting” was introduced in (Adler et al., 2011) to describe the restoration of missing or corrupted audio segments, although related notions had appeared under names such as audio interpolation (Marafioti et al., 2020), audio extrapolation (Lieb and Stark, 2018), missing sample reconstruction, waveform substitution (Adler et al., 2011), and imputation (Lieb and Stark, 2018).
Classical methods mainly targeted short gaps (<100<100 ms), often assuming local stationarity. Auto-regressive models (Adler et al., 2011) predicted missing samples from past ones but degrade on longer gaps. Sparse reconstruction in time–frequency domains like STFT or Gabor (Lieb and Stark, 2018; Adler et al., 2011) exploited algorithms such as OMP (Tropp and Gilbert, 2007), later improved by adaptive dictionaries (e.g., learned Gabor atoms (Tauböck et al., 2020)). NMF-based methods (Mokrỳ et al., 2023) filled spectrogram gaps via low-rank factorization, with probabilistic variants improving robustness. Other structured priors include sinusoidal modeling for harmonic signals (Adler et al., 2011) and graph-based regularization: (Perraudin et al., 2018) used a graph Laplacian on self-similarity matrices to borrow content from uncorrupted regions.
While effective for very short gaps, these models fail for longer ones due to violated assumptions of stationarity or sparsity. Deep learning approaches were introduced to overcome this challenge. CNNs first inpainted spectrograms, with (Marafioti et al., 2019) using a U-Net style context encoder for tens-of-milliseconds gaps. GANs further improved realism: (Ebner and Eltelt, 2020) proposed a Wasserstein GAN with multi-context conditioning for gaps up to 500 ms, while (Marafioti et al., 2020) extended this with GACELA, a multi-scale GAN handling 1–1.5 s gaps. These data-driven models substantially outperform classical approaches, especially on long gaps with complex temporal or musical structure. More recently, A score-aware GAN framework (Aironi et al., 2023a) further improved music inpainting by incorporating symbolic losses, outperforming GACELA on medium gaps.
@@ 2.2 Audio Inpainting using Diffusion Models
Diffusion models, or score-based generative models, have recently transformed vision and audio generation. They define a forward noising process and learn its reversal (e.g., DDPMs (Ho et al., 2020), score-SDEs (Song et al., 2020)). In audio, they support waveform and spectrogram generation conditioned on context. For instance, DiffWave (Kong et al., 2020) synthesizes raw waveforms for vocoding, while other works apply diffusion in latent or time–frequency domains (STFT, CQT) to exploit structure (Moliner et al., 2023; Moliner and Välimäki, 2023). Most use continuous diffusion, though discrete approaches exist: DiffSound (Yang et al., 2023) applies discrete diffusion to quantized spectrogram tokens. Conceptually, continuous diffusion perturbs real values, yielding smooth interpolations, whereas discrete diffusion masks/replaces categorical tokens, enabling multimodal reconstructions but with possible quantization artifacts (Ho et al., 2020; Song et al., 2020; Yang et al., 2023).
Diffusion has recently been adapted to inpainting. (Moliner and Välimäki, 2023) propose a zero-shot unconditional model on CQT spectrograms, with posterior sampling to enforce data consistency (Kawar et al., 2022; Chung et al., 2022). This handles arbitrary gaps, matching baselines on short ones and surpassing them up to 300 ms. Similarly, CQT-Diff (Moliner et al., 2023) uses an invertible CQT front-end for inpainting, declipping, and bandwidth extension, achieving state-of-the-art restoration. MAID (Liu et al., 2023), instead, trains a conditional diffusion model for music inpainting guided by context and auxiliary cues (e.g., pitch), enabling longer gaps but requiring task-specific supervision. (Aironi et al., 2023a) further propose score-aware conditioning, aligning the score function with masked regions for musically coherent results.
Architectures differ in operating domain (raw waveform vs. spectrogram/latent space) and conditioning. Zero-shot approaches (Moliner and Välimäki, 2023) enforce context consistency during inference, while conditional models integrate masked audio or side information during training. Masking usually involves contiguous spans: Moliner et al. test gaps from 25-300 ms, while GAN baselines address seconds-long gaps (Marafioti et al., 2020; Ebner and Eltelt, 2020). Diffusion models degrade gracefully with gap length: all interpolate short gaps well, but diffusion’s generative prior yields more plausible completions beyond 100 ms compared to sparsity- or AR-based methods.
In summary, diffusion-based inpainting combines strong generative priors with flexible conditioning. Open questions remain on continuous vs. discrete diffusion (Yang et al., 2023) and conditioning strategies (posterior sampling vs. trained conditional networks).
@@ 3.1 Discrete Diffusion Models
Continuous Diffusion Models (CDMs) (Ho et al., 2020; Song et al., 2020) have achieved remarkable success across various generative modeling tasks, particularly in the image domain (Dhariwal and Nichol, 2021). The high-dimensional nature of raw audio data adds to the difficulty of directly modeling it within the diffusion framework. An effective strategy to address this issue is to first compress continuous audio signals into compact, discrete representations, such as quantized tokens derived from a learned codebook (Défossez et al., 2022). DDMs, which operate in a token space, have recently demonstrated strong performance in the domain of natural language generation (Lou et al., 2024; Nie et al., 2025; Sahoo et al., 2024). Similarly to CDM, DDMs comprise a forward noising process that progressively corrupts an initial sample $x_{0}$ into a maximally noisy (fully masked) version $x_{T}$, and a reverse denoising process trained to reconstruct $x_{0}$ from $x_{T}$.
@@
Existing work (Lou et al., 2024) shows that the forward diffusion in discrete space can be formulated using a token-level transition rate matrix $Q_{\text{tok}}\in\mathbb{R}^{n\times n}$, where $n$ is the vocabulary size. The forward transition probability of a single token $x_{0}$ at time $t$ is then given by
@@
The reverse process is described by a new reverse transition matrix (Sun et al., 2022; Kelly, 1981) which is defined as $\bar{Q}_{t}(y,x)=\tfrac{p_{t}(y)}{p_{t}(x)}Q_{t}(x,y)$ and $\bar{Q}_{t}(x,x)=-\sum_{y\neq x}\bar{Q}_{t}(y,x)$.
To simulate the reverse process, it is sufficient to estimate the ratio $\frac{p_{t}(y)}{p_{t}(x)}$, referred to as the concrete-score (Lou et al., 2024; Meng et al., 2022), which can be approximated by a neural score function: $s_{\theta}(x,t)\approx\left[\tfrac{p_{t}(y)}{p_{t}(x)}\right]_{y\in\mathcal{X},\,y\neq x}$.
@@
To train this score function, recent work has proposed the Diffusion Weighted Denoising Score Entropy (DWDSE) objective (Lou et al., 2024), which is defined as:
@@ 4 Audio Inpainting using Discrete Diffusion Model
To the best of our knowledge, our method Audio Inpainting using Discrete Diffusion (AIDD) is the first to leverage a discrete diffusion model for audio inpainting. AIDD comprises three main components: (1) audio tokenization using WavTokenizer (Ji et al., 2024), (2) generative modeling using a DDM (Lou et al., 2024), and (3) waveform reconstruction via token decoding (Ji et al., 2024). Furthermore, we propose a new training derivative-based loss, and a span-based masking method, extending the approach of (Lou et al., 2024) to more effectively address the audio inpainting task.
@@ 4.1 The Proposed Architecture
AIDD consists of WavTokenizer (Ji et al., 2024) that transforms high-dimensional raw audio signals into compact sequences of discrete tokens and reconstructs it back to raw audio, and a Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023) which learns to predict masked tokens through the reverse diffusion process. An overview of the proposed framework is illustrated in Figure 1.
Figure 1: Our method operates on audio signals with missing (silent) segments. During inference, the input waveform, containing a single or multiple silence gaps is processed by the WavTokenizer encoder, which converts the audio into a discrete sequence of tokens. Next, a DiT performs inpainting by iteratively predicting the masked tokens, resulting in reconstructed token sequence. Finally, the reconstructed tokens are passed through the WavTokenizer’s decoder to synthesize the output audio waveform in the masked part. During training, token sequences are corrupted with span-based masking at randomly sampled timesteps, and the DiT is optimized to predict the concrete score using the DWDSE objective, complemented by the derivative-based loss.
@@
The first step in our method is to convert the raw audio waveform into discrete audio tokens that can be efficiently processed and trained by the DDM. We utilize a pre-trained audio tokenizer WavTokenizer (Ji et al., 2024), which compresses high-resolution audio into a compact sequence of discrete tokens using a single quantizer. Despite its extreme compression, WavTokenizer preserves high reconstruction fidelity and rich semantic content. Unlike conventional audio inpainting frameworks that operate directly on the audio waveform or the spectrogram (Moliner and Välimäki, 2023), audio tokenization allows us to formulate audio inpainting as a discrete sequence completion task. We also evaluated an alternative tokenizer, UniCodec (Jiang et al., 2025), which is based on a transformer (Vaswani et al., 2017) architecture. Results are presented in Section 7.
@@
At the core of our method is a DiT architecture (Peebles and Xie, 2023), which integrates time conditioning into a standard encoder-only transformer (Vaswani et al., 2017; Devlin et al., 2019) and, following the work of (Lou et al., 2024), we incorporate rotary positional encoding (Su et al., 2024). To operate effectively in the discrete token space, we use the denoising score entropy formulation, guiding the reverse diffusion dynamics as seen in Eq. 3.